geometry problems:
I'm uncertain as how to set these up.

the length of a rectangle is two feet more than twice the width.
the area is 144ft^2. find the length and the width of the rectangle.

Answer by venugopalramana(3286)

the length=L SAY of a rectangle is two feet more than twice the width=W SAY.
the area is 144ft^2. find the length and the width of the rectangle.
TWICE WIDTH =2W
2 FEET MORE =2W+2=2(W+1)
L=2W+2
AREA=A=L*W=2(W+1)W=144
W(W+1)=144/2=72
W^2+W-72=0
W^2+9W-8W-72=0
W(W+9)-8(W+9)=0
(W-8)(W+9)=0
W=8
L=2(8+1)=18